find the value of a
A line has slope -5/2 and passes through points (8,-4) and (a,6)

Answer by TimothyLamb(4379)

Get the equation in point-slope form: y - y1 = m(x - x1)
---
using:
m = (-5/2)
p1 = (8, -4)
p2 = (a, 6)
---
y - y1 = m(x - x1)
6 - -4 = (-5/2)(a - 8)
10 = (-5/2)a + (5/2)8
10 = (-5/2)a + 20
(5/2)a = 20 - 10
(5/2)a = 10
a = 10(2/5)
---
Answer:
a = 4
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php